# TACA: Rethinking Cross-Modal Interaction in Multimodal Diffusion Transformers

### v2

@@ Abstract
Multimodal Diffusion Transformers (MM-DiTs) have achieved remarkable progress in text-driven visual generation. However, even state-of-the-art MM-DiT models like FLUX struggle with achieving precise alignment between text prompts and generated content. We identify two key issues in the attention mechanism of MM-DiT, namely 1) the suppression of cross-modal attention due to token imbalance between visual and textual modalities and 2) the lack of timestep-aware attention weighting, which hinder the alignment. To address these issues, we propose Temperature-Adjusted Cross-modal Attention (TACA), a parameter-efficient method that dynamically rebalances multimodal interactions through temperature scaling and timestep-dependent adjustment. When combined with LoRA fine-tuning, TACA significantly enhances text-image alignment on the T2I-CompBench benchmark with minimal computational overhead. We tested TACA on state-of-the-art models like FLUX and SD3.5, demonstrating its ability to improve image-text alignment in terms of object appearance, attribute binding, and spatial relationships. Our findings highlight the importance of balancing cross-modal attention in improving semantic fidelity in text-to-image diffusion models. Our codes are publicly available at https://github.com/Vchitect/TACA.
@@
footnotetext: *Equal Contribution. Project Leader. Corresponding Author.
@@ 1 Introduction
Diffusion models [13, 39], driven by iterative denoising processes, have emerged as a powerful paradigm in generative modeling. The field has witnessed significant architectural evolution, beginning with U-Net-based designs [37] that dominated early diffusion models [9, 36, 32, 27]. Recent advances introduced transformer-based architectures through Diffusion Transformers (DiT) [29, 5], demonstrating superior scalability and training stability. This progression culminated in Multimodal Diffusion Transformers (MM-DiT) [10], which unify text and visual tokens through a concatenated self-attention mechanism, resulting in state-of-the-art text-to-image/video models like Stable Diffusion 3/3.5 [10, 40], FLUX [20], CogVideo [14, 48], and HunyuanVideo [19].
@@ 2.2 Text-to-Image Alignment
Prior research has explored generating images from text prompts using pre-trained models without requiring further training, employing techniques such as CLIP-guided optimization [25, 12] to align images with text by optimizing CLIP scores within the model’s latent space, and cross-attention-based approaches [7] to enhance spatial layout and details in generated images, thereby improving adherence to the textual description’s structure.
@@ 3.2 Suppression of Cross-Attention and Timestep-Insensitive Weighting in MM-DiT
Suppression of Cross-Attention in Unified Softmax. This issue stems from the numerical asymmetry between the number of visual and text tokens ($N_{x}\gg N_{c}$), which creates a systematic bias in attention weight distribution. Consider the attention computation for visual tokens in Eq. 4. Each visual token’s attention weights over text tokens ($\boldsymbol{Q}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{txt}}^{T}$) must compete against visual-visual interactions ($\boldsymbol{Q}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{vis}}^{T}$) in the softmax denominator. Formally, the probability of the $i$-th visual token attending to the $j$-th text token guidance becomes:
@@
where $s_{ik}^{\mathrm{vt}}$ $=$ $\boldsymbol{Q}^{(i)}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{txt}}^{T\,(k)}/\sqrt{D}$ and $s_{ik}^{\mathrm{vv}}$ $=$ $\boldsymbol{Q}^{(i)}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{vis}}^{T\,(k)}/\sqrt{D}$. When $N_{\mathrm{vis}}\gg N_{\mathrm{txt}}$, the exponential sum over visual-visual interactions dominates the denominator, even if individual $s_{ik}^{\mathrm{vv}}$ values are modest. For example, when using FLUX.1 Dev [20] to generate a 1024 $\times$ 1024 image, we have $N_{\mathrm{vis}}/N_{\mathrm{txt}}=4096/512=8$. In this case, the visual-text cross-attention probabilities would be much lower than in the typical paradigm:
@@
Timestep-Insensitive QK Weighting. MM-DiT’s current architecture employs timestep-agnostic projection of latent states into query and key vectors. This approach fails to account for the evolving influence of textual guidance throughout the denoising process. As illustrated in Fig. 4, the initial denoising steps are crucial for establishing the image’s global layout, heavily influenced by the text prompt. Consequently, the cross-attention mechanism, responsible for integrating textual information, should be weighted more heavily than visual self-attention during these early stages. MM-DiT’s static weighting strategy, therefore, limits its ability in optimally leveraging textual guidance and adapting to the changing demands of the denoising process.
Formally, when $t$ is large (i.e., early in the denoising process) and cross-modal guidance should dominate, $s_{ik}^{\mathrm{vt}}$ values fail to receive proportionally larger magnitudes compared to $s_{ik}^{\mathrm{vv}}$. Since $\boldsymbol{W}^{Q}$ and $\boldsymbol{W}^{K}$ are optimized for global performance across all timesteps, they cannot focus on amplifying text-visual interactions in the early stages. This potentially leads to the overall layout of the generated image not aligning with the text prompt.
Figure 5: Temperature scaling helps visual-text alignment. From this figure, we can see that as the temperature scaling factor $\gamma$ increases, the characteristics of “brown backpack”, “glass mirror” and “black stomach” become more obvious.
@@ 3.3 Temperature-Adjusted Cross-modal Attention
To address the two issues mentioned in Section 3.2, we propose Temperature-Adjusted Cross-modal Attention (TACA), a simple yet effective modification to the attention mechanism of MM-DiT. Our approach introduces two key innovations, namely modality-specific temperature scaling and timestep-dependent adjustment of cross-modal interactions.
Modality-Specific Temperature Scaling. To mitigate the suppression of cross-attention caused by the dominance of visual tokens ($N_{\mathrm{vis}}\gg N_{\mathrm{txt}}$), we amplify the logits of visual-text interactions through a temperature coefficient $\gamma>1$. The modified attention probability for visual-text interaction becomes:
@@
This scaling effectively rebalances the competition in softmax by increasing the relative weight of cross-modal interactions. The $\gamma$ coefficient acts as a signal booster for text-guided attention. As shown in Fig. 5, the generated image and text prompt become more consistent as $\gamma$ increases.
Timestep-Dependent Adjustment. To compensate for the insensitivity of QK weights with respect to the timestep, we make $\gamma$ timestep-dependent to account for the varying importance of cross-attention during denoising based on the observations in Fig. 4. Specifically, we employ a piecewise function:
@@
where $t_{\mathrm{thresh}}$ is a threshold for the timestep that separates the layout formation and detail refinement phases. This design aligns with the denoising dynamics where early steps (i.e., large $t$) require strong text guidance to establish image composition and later steps (i.e., small $t$) focus on visual details when self-attention dominates.
@@
LoRA Training for Artifact Suppression. While temperature scaling in TACA significantly improves visual-image alignment, the amplified cross-modal attention logits can alter the output distribution of the denoising process, occasionally introducing artifacts such as distorted object boundaries or inconsistent textures. To mitigate this, we employ Low-Rank Adaptation (LoRA) [15] to fine-tune the model, encouraging it to recover the real image distribution while preserving the benefits of temperature scaling.
@@
Figure 7: Comparison of samples generated by FLUX.1 Dev and Stable Diffusion 3.5 Medium with and without TACA. Our method enhances the semantic representation of the primary object within the text prompt, improves spatial relationships, and maintains image quality without degradation.
Table 1: Comparison of alignment evaluation on T2I-CompBench [16] for FLUX.1-Dev-based and SD3.5-Medium-based models. The best results for each model group are highlighted in $\mathbf{bold}$.
@@ 4.1 Experiment Settings
Evaluation Metrics and Datasets. We evaluate our method on the T2I-CompBench benchmark [16], a comprehensive evaluation suite for text-to-image alignment. All experiments use the LAION dataset [38] with captions refined by the LLaVA model [24] to enhance semantic precision. We randomly sampled 10K image-text pairs as the training dataset for our LoRA model. To ensure reproducibility throughout all evaluation phases, the random seed is fixed to 42, while all other parameters remain at their default values as provided by the Diffusers library [43].
Implementation Details. We conduct experiments on a single NVIDIA A100 80GB GPU using the ai-toolkit codebase [28], with LoRA adapters implemented for FLUX.1 Dev [20] and SD3.5 Medium [40] models. We adopt the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a batch size of $4$ for training. We evaluate two LoRA configurations: $(r,\alpha)$ = $(16,16)$ and $(64,64)$.
To emphasize semantic alignment, we sample timesteps $t\geq t_{\mathrm{thresh}}=970$ within the range $t\in(0,1000)$. In the Diffusers’ flowmatching scheduler, a $30$-step denoising process allocates the first three steps to $t\in(970,1000]$ (i.e., the initial $10\%$ of the diffusion process), while the remaining $27$ steps cover $t\in[0,970)$. Setting $t_{\mathrm{thresh}}=970$ focuses training on these early steps where semantic information is most prominent.
@@ 4.2 Main Results
Quantitative Comparison. To quantitatively evaluate the effectiveness of our proposed TACA, we conduct a comprehensive comparison against baseline models. Table 1 presents the alignment performance of FLUX.1-Dev and SD3.5-Medium models, respectively, with and without the integration of TACA. For FLUX.1-Dev, the incorporation of TACA, particularly with a rank $r=64$, consistently improves performance across all Attribute Binding metrics and Spatial Relationship. Similarly, for SD3.5-Medium, TACA with $r=64$ yields significant gains in Attribute Binding and Spatial Relationship, and TACA with $r=16$ achieves the best performance on Non-Spatial Relationship and Complex prompt evaluation. These results demonstrate that TACA effectively enhances the alignment capabilities of different MM-DiT models across various dimensions of text-to-image generation quality.
Image Quality Evaluation. We use widely adopted image quality assessment models MUSIQ [17] and MANIQA [47] to evaluate visual quality. As shown in the Table 2, TACA improves text-image alignment without sacrificing image quality on both SD3.5 and FLUX.
@@
User Study. We invited 50 participants for our user study. From the T2I-Compbench [16] dataset, we sampled 25 prompts and generated images using FLUX.1 Dev model with and without our TACA method. These images, along with their corresponding text prompts, were presented to the participants. Participants were asked to indicate their preferred image based on three criteria, namely overall visual appeal, attribute (color/shape/texture) quality, and text-image alignment. The results, as summarized in Table 3, demonstrate that a majority of participants favored the images generated by the model incorporating the TACA method. This suggests that our method yields improvements in text alignment and does not ruin image quality.
@@ 4.3 Ablation Study
Figure 8: The effect of LoRA training. (a) Qualitative results comparing training-free image generation to generation with LoRA. “Training free’ examples show artifacts such as a floating bowl and distorted structures, which are significantly reduced by LoRA training. (b) Quantitative evaluation of CLIP Scores for training-free and LoRA-trained models across the denoising steps, showing that LoRA maintains the strong visual-text alignment and does not detract from the semantic benefits of our approach.
Effect of Temperature Scaling Factor ($\gamma_{0}$). The temperature scaling factor $\gamma_{0}$ plays a crucial role in modulating the influence of textual guidance during the denoising process. We explore the impact of different $\gamma_{0}$ values on compositional generation performance. Table 4 presents the results.
@@
To understand the mechanism behind this improvement, we analyze the CLIP similarity between the predicted intermediate latent representations and the text prompt at each denoising step for varying $\gamma_{0}$. As shown in Figure 8 (b), increasing $\gamma_{0}$ leads to a higher CLIP similarity, particularly in the initial denoising steps. This indicates that TACA effectively enhances the visual-text alignment early in the generation process, guiding the model towards generating images that are more consistent with the textual description.
Sensitivity of $\gamma_{0}$ and $t_{\text{thresh}}$. We further investigate the sensitivity of our method to the choice of $\gamma_{0}$ and the threshold timestep $t_{\text{thresh}}$ beyond which TACA is applied. Table 5 presents the average Attribute (Color, Shape, Texture) and Spatial scores for different values of $\gamma_{0}$ and $t_{\text{thresh}}$ on both FLUX and SD3.5 models.
Table 5: Sensitivity analysis of $\gamma_{0}$ and $t_{\text{thresh}}$. We evaluate the average Attribute (mean score of Color, Texture and Shape) and Spatial scores for different values of $\gamma_{0}$ and $t_{\text{thresh}}$ on FLUX.1-Dev and SD3.5-Medium models. The baseline corresponds to the respective models without TACA.
@@
Attribute Spatial Attribute Spatial Attribute Spatial Attribute Spatial
@@
As shown in Table 5, the performance exhibits minimal variation across a reasonable range of both $\gamma_{0}$ (e.g., 1.15 to 1.25) and $t_{\text{thresh}}$ (e.g., 930 to 970). This indicates that our method is not overly sensitive to the precise selection of these parameters, suggesting practical robustness. Furthermore, this robustness is observed across different base models=, highlighting the general applicability of the tested parameter ranges.
The Effect of LoRA Training. In the original TACA method, the introduction of factor $\gamma(t)$ induces a shift in the output distribution of each attention layer. These modified outputs are subsequently processed by the feed-forward networks within the transformer blocks. Consequently, the overall output distribution of the diffusion transformer deviates from the distribution inherent in real images, which manifests as visual artifacts like unsupported floating bowls and distorted bridge connections in Fig 8 (a). To address this issue, we hypothesize that training a LoRA [15] module can effectively mitigate these artifacts. The rationale is that by fine-tuning the attention layer weights with a limited number of training samples, the LoRA module enables the modified model to readjust its output distribution to better align with the real image distribution.
Empirical findings from our experiments demonstrate that the incorporation of LoRA significantly enhances image quality and effectively mitigates these unrealistic artifacts, as evidenced in Fig. 8 (a). Concurrently, we evaluated whether the introduction of LoRA compromises the semantic enhancement facilitated by the temperature coefficient $\gamma(t)$. The comparative analysis of CLIP Scores for training-free and LoRA configurations for 50 samples, presented in Fig. 8 (b), reveals that LoRA exerts a negligible impact on visual-text alignment.

### v3

# Rethinking Cross-Modal Interaction in Multimodal Diffusion Transformers

@@ Abstract
Multimodal Diffusion Transformers (MM-DiTs) have achieved remarkable progress in text-driven visual generation. However, even state-of-the-art MM-DiT models like FLUX struggle with achieving precise alignment between text prompts and generated content. We identify two key issues in the attention mechanism of MM-DiT, namely 1) the suppression of cross-modal attention due to token imbalance between visual and textual modalities and 2) the lack of timestep-aware attention weighting, which hinder the alignment. To address these issues, we propose Temperature-Adjusted Cross-modal Attention (TACA), a parameter-efficient method that dynamically rebalances multimodal interactions through temperature scaling and timestep-dependent adjustment. When combined with LoRA fine-tuning, TACA significantly enhances text-image alignment on the T2I-CompBench benchmark with minimal computational overhead. We tested TACA on state-of-the-art models like FLUX and SD3.5, demonstrating its ability to improve text-image alignment in terms of object appearance, attribute binding, and spatial relationships. Our findings highlight the importance of balancing cross-modal attention in improving semantic fidelity in text-to-image diffusion models. Our codes are publicly available at https://github.com/Vchitect/TACA.
@@
footnotetext: *Equal Contribution. Corresponding Author.
@@ 1 Introduction
Diffusion models [13, 40], driven by iterative denoising processes, have emerged as a powerful paradigm in generative modeling and various visual generation tasks [46, 26, 47, 52]. The field has witnessed significant architectural evolution, beginning with U-Net-based designs [38] that dominated early diffusion models [9, 37, 33, 28]. Recent advances introduced transformer-based architectures through Diffusion Transformers (DiT) [30, 5], demonstrating superior scalability and training stability. This progression culminated in Multimodal Diffusion Transformers (MM-DiT) [10], which unify text and visual tokens through a concatenated self-attention mechanism, resulting in state-of-the-art text-to-image/video models like Stable Diffusion 3/3.5 [10, 41], FLUX [20], CogVideo [14, 51], and HunyuanVideo [19].
@@ 2.2 Text-to-Image Alignment
Prior research has explored generating images from text prompts using pre-trained models without requiring further training. Some employ techniques such as CLIP-guided optimization [25, 12] to align images with text by optimizing CLIP scores within the model’s latent space. Additionally, cross-attention-based approaches [7] are used to enhance spatial layout and details in generated images, thereby improve adherence to the textual description’s structure.
@@ 3.2 Suppression of Cross-Attention and Timestep-Insensitive Weighting in MM-DiT
Suppression of Cross-Attention This issue stems from the numerical asymmetry between the number of visual and text tokens ($N_{x}\gg N_{c}$), which creates a systematic bias in attention weight distribution. Consider the attention computation for visual tokens in Eq. 6. Each visual token’s attention weights over text tokens ($\boldsymbol{Q}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{txt}}^{T}$) must compete against visual-visual interactions ($\boldsymbol{Q}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{vis}}^{T}$) in the softmax denominator. Formally, the probability of the $i$-th visual token attending to the $j$-th text token guidance becomes:
@@
where $s_{ik}^{\mathrm{vt}}$ $=$ $\boldsymbol{Q}^{(i)}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{txt}}^{T\,(k)}/\sqrt{D}$ and $s_{ik}^{\mathrm{vv}}$ $=$ $\boldsymbol{Q}^{(i)}_{\mathrm{vis}}\boldsymbol{K}_{\mathrm{vis}}^{T\,(k)}/\sqrt{D}$. When $N_{\mathrm{vis}}\gg N_{\mathrm{txt}}$, the sum over visual-visual interactions dominates the denominator, even if individual $s_{ik}^{\mathrm{vv}}$ values are modest. For example, when using FLUX.1 Dev [20] to generate a 1024 $\times$ 1024 image, we have $N_{\mathrm{vis}}/N_{\mathrm{txt}}=4096/512=8$. In this case, the visual-text cross-attention probabilities would be much lower than in the typical paradigm:
@@
Timestep-Insensitive QK Weighting MM-DiT’s current architecture employs timestep-agnostic projection of latent states into query and key vectors. This approach fails to account for the evolving influence of textual guidance throughout the denoising process. As illustrated in Fig. 4, the initial denoising steps are crucial for establishing the image’s global layout, heavily influenced by the text prompt. Consequently, the cross-attention mechanism, responsible for integrating textual information, should be weighted more heavily than visual self-attention during these early stages. MM-DiT’s static weighting strategy, therefore, limits its ability in optimally leveraging textual guidance and adapting to the changing demands of the denoising process.
Formally, when $t$ is large (i.e., early in the denoising process) and cross-modal guidance should dominate, $s_{ik}^{\mathrm{vt}}$ values fail to receive proportionally larger magnitudes compared to $s_{ik}^{\mathrm{vv}}$. Since $\boldsymbol{W}^{Q}$ and $\boldsymbol{W}^{K}$ are optimized for global performance across all timesteps, they cannot focus on amplifying visual-text interactions in the early stages. This potentially leads to the overall layout of the generated image not aligning with the text prompt.
Figure 5: Temperature scaling helps text-image alignment. From this figure, we can see that as the temperature scaling factor $\gamma$ increases, the characteristics of “brown backpack”, “glass mirror” and “black stomach” become more obvious.
@@ 3.3 Temperature-Adjusted Cross-modal Attention
To address the issues mentioned in Section 3.2, we propose Temperature-Adjusted Cross-modal Attention (TACA), a simple yet effective modification to the attention mechanism of MM-DiT. Our approach introduces two key innovations, namely modality-specific temperature scaling and timestep-dependent adjustment of cross-modal interactions.
Modality-Specific Temperature Scaling To mitigate the suppression of cross-attention caused by the dominance of visual tokens ($N_{\mathrm{vis}}\gg N_{\mathrm{txt}}$), we amplify the logits of visual-text interactions through a temperature coefficient $\gamma>1$. The modified attention probability for visual-text interaction becomes:
@@
This scaling effectively rebalances the competition in softmax by increasing the relative weights of cross-modal interactions. The $\gamma$ coefficient acts as a signal booster for text-guided attention. As shown in Fig. 5, the generated image and text prompt become more consistent as $\gamma$ increases.
Timestep-Dependent Adjustment To compensate for the insensitivity of QK weights with respect to the timestep, we make $\gamma$ timestep-dependent to account for the varying importance of cross-attention during denoising based on the observations in Fig. 4. Specifically, we employ a piecewise function:
@@
where $t_{\mathrm{thresh}}$ is a threshold for the timestep that separates the layout formation and detail refinement phases. This design aligns with the denoising dynamics where early steps (i.e., large $t$) require strong text guidance to establish image composition and later steps (i.e., small $t$) focus on visual details when self-attention dominates. By effecting the attention map, image tokens can better attend to the relevant text tokens, as shown in Fig. 6.
@@
LoRA Training for Artifact Suppression While temperature scaling in TACA significantly improves text-image alignment, the amplified cross-modal attention logits can alter the output distribution of the denoising process, occasionally introducing artifacts such as distorted object boundaries or inconsistent textures. To mitigate this, we employ Low-Rank Adaptation (LoRA) [15] to fine-tune the model, encouraging it to recover the real image distribution while preserving the benefits of temperature scaling.
@@
Figure 7: Comparison of samples generated by FLUX.1 Dev and Stable Diffusion 3.5 Medium with and without TACA.
Table 1: Comparison of alignment evaluation on T2I-CompBench [16] for FLUX.1-Dev-based and SD3.5-Medium-based models. The best results for each model group are highlighted in bold.
@@ 4.1 Experiment Settings
Evaluation Metrics and Datasets We evaluate our method on the T2I-CompBench benchmark [16], a comprehensive evaluation suite for text-to-image alignment. All experiments use the LAION dataset [39] with captions refined by the LLaVA model [24] to enhance semantic precision. We randomly sampled 10K image-text pairs as the training dataset for our LoRA model. To ensure reproducibility throughout all evaluation phases, the random seed is fixed to 42, while all other parameters remain at their default values as provided by the Diffusers library [44].
Implementation Details We conduct experiments on a single NVIDIA A100 80GB GPU using the ai-toolkit codebase [29], with LoRA adapters implemented for FLUX.1 Dev [20] and SD3.5 Medium [41] models. We adopt the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a batch size of $4$ for training. We evaluate two LoRA configurations: $(r,\alpha)$ = $(16,16)$ and $(64,64)$.
To emphasize semantic alignment, we sample timesteps $t\geq t_{\mathrm{thresh}}=970$ within the range $t\in(0,1000)$. In the flow matching scheduler, a $30$-step denoising process allocates the first three steps to $t\in(970,1000]$ (i.e., the initial $10\%$ of the diffusion process), while the remaining $27$ steps cover $t\in[0,970)$. Setting $t_{\mathrm{thresh}}=970$ focuses training on these early steps where semantic information is most prominent.
@@ 4.2 Main Results
Quantitative Comparison  To quantitatively evaluate the effectiveness of our proposed TACA, we conduct a comprehensive comparison against baseline models. Table 1 presents the alignment performance of FLUX.1-Dev and SD3.5-Medium models, respectively, with and without the integration of TACA. For FLUX.1-Dev, the incorporation of TACA, particularly with a rank $r=64$, consistently improves performance across all Attribute Binding metrics and Spatial Relationship. Similarly, for SD3.5-Medium, TACA with $r=64$ yields significant gains in Attribute Binding and Spatial Relationship, and TACA with $r=16$ achieves the best performance on Non-Spatial Relationship and Complex prompt evaluation. These results demonstrate that TACA effectively enhances the alignment capabilities of different MM-DiT models across various dimensions of text-to-image generation quality.
Image Quality Evaluation We use widely adopted image quality assessment models MUSIQ [17] and MANIQA [50] to evaluate visual quality. As shown in the Table 2, TACA improves text-image alignment without sacrificing image quality on both SD3.5 and FLUX. Additionally, Fig. 7 presents further visual comparison results.
@@
User Study We invited 50 participants for our user study. From the T2I-Compbench [16] dataset, we sampled 25 prompts and generated images using FLUX.1 Dev model with and without our TACA method. These images, along with their corresponding text prompts, were presented to the participants. Participants were asked to indicate their preferred image based on three criteria, namely overall visual appeal, attribute (color/shape/texture) quality, and text-image alignment. The results, as summarized in Table 3, demonstrate that a majority of participants favored the images generated by the model incorporating the TACA method. This suggests that our method yields improvements in text alignment and does not ruin image quality.
@@ 4.3 Ablation Study
Figure 8: The effect of LoRA training. (a) On the left, we show qualitative results comparing training-free image generation to generation with LoRA. The ‘training-free’ examples exhibit artifacts, such as the floating bowl, which are significantly reduced by LoRA training. (b) On the right, we present a quantitative evaluation of CLIP Scores for training-free and LoRA-trained models across the denoising steps, demonstrating that LoRA maintains strong text-image alignment and does not detract from the semantic benefits of our approach.
Effect of Temperature Scaling Factor ($\gamma_{0}$) The temperature scaling factor $\gamma_{0}$ plays a crucial role in modulating the influence of textual guidance during the denoising process. We explore the impact of different $\gamma_{0}$ values on compositional generation performance. Table 4 presents the results.
@@
To understand the mechanism behind this improvement, we analyze the CLIP similarity between the predicted intermediate latent representations and the text prompt at each denoising step for varying $\gamma_{0}$. As shown in Figure 8 (b), increasing $\gamma_{0}$ leads to a higher CLIP similarity, particularly in the initial denoising steps. This indicates that TACA effectively enhances the text-image alignment early in the generation process, guiding the model towards generating images that are more consistent with the textual description.
Sensitivity of $\gamma_{0}$ and $t_{\text{thresh}}$e further investigate the sensitivity of our method to the choice of $\gamma_{0}$ and the threshold timestep $t_{\text{thresh}}$ beyond which TACA is applied. Table 5 presents the average Attribute (Color, Shape, Texture) and Spatial scores for different values of $\gamma_{0}$ and $t_{\text{thresh}}$ on both FLUX and SD3.5 models.
Table 5: Sensitivity analysis of $\gamma_{0}$ and $t_{\text{thresh}}$. The baseline corresponds to the respective models without TACA.
@@
attr spa attr spa attr spa attr spa
@@
As shown in Table 5, the performance exhibits minimal variation across a reasonable range of both $\gamma_{0}$ (e.g., 1.15 to 1.25) and $t_{\text{thresh}}$ (e.g., 930 to 970). This indicates that our method is not overly sensitive to the precise selection of these parameters, suggesting practical robustness. Furthermore, this robustness is observed across different base models, highlighting the general applicability of the tested parameter ranges.
The Effect of LoRA Training In the original TACA method, the introduction of factor $\gamma(t)$ induces a shift in the output distribution of each attention layer. These modified outputs are subsequently processed by the feed-forward networks within the transformer blocks. Consequently, the overall output distribution of the diffusion transformer deviates from the distribution inherent in real images, which manifests as visual artifacts like unsupported floating bowls and distorted bridge connections in Fig 8 (a). To address this issue, we hypothesize that training a LoRA [15] module can effectively mitigate these artifacts. The rationale is that by fine-tuning the attention layer weights with a limited number of training samples, the LoRA module enables the modified model to readjust its output distribution to better align with the real image distribution.
Empirical findings from our experiments demonstrate that the incorporation of LoRA significantly enhances image quality and effectively mitigates these unrealistic artifacts, as evidenced in Fig. 8 (a). Concurrently, we evaluated whether the introduction of LoRA compromises the semantic enhancement facilitated by the temperature coefficient $\gamma(t)$. The comparative analysis of CLIP Scores for training-free and LoRA configurations for 50 samples, presented in Fig. 8 (b), reveals that LoRA exerts a negligible impact on text-image alignment.